Zoron is 10 years older than his sister.In 3 years,he will be twice as old as she will be then. How old are they now?
Answer by stanbon(48546)

Zoron is 10 years older than his sister.In 3 years,he will be twice as old as she will be then. How old are they now?
---------------------------------------------
NOW DATA
Let his sister's age be "x" yrs old
Then Zoron is "x+10" yrs old.
----------------------------------------
3 yrs from now DATA:
His sister will be "x+3" yrs old
Zoron will be "x+13" yrs old
------------------------------
Equations:
x+13 = 2(x+3)
x + 13 = 2x + 6
x = 7 (his sister is now 7 yrs old
Zonron is now x+10 = 17 yrs old
====================================
Cheers,
Stan H.